What is the force of the gravity that is exerted on the Earth by the sun? The mass of the sun and earth are 1.99 × 10^30kg and 5.97×10^24kg respectively. The distance between the earth and tye sun is 1.496 × 10 ^11 m.

1
Expert's answer
2022-03-16T09:41:42-0400

The force of the gravity

"F=G\frac{m_1m_2}{r^2}"

"F=6.67*10^{-11}*\frac{1.99*10^{30}*5.97*10^{24}}{(1.496*10^{11})^2}=3.36*10^{22}\:\rm N"